OpenStudy (anonymous):

How do I determine the domain and range of this function F(x) = -3x+8

OpenStudy (anonymous):

Domian = All real numbers Since there is no restriction

OpenStudy (anonymous):

Domain=R Range=R Because no interval in this question

OpenStudy (anonymous):

as the function is linear its domain is all real numbers

OpenStudy (anonymous):

For Range take y = f(x) y = -3x +8 8 - y = 3x x = (8-y)/3 So.....Range = All REAL NUMBERS

OpenStudy (anonymous):

range is also real numbers

OpenStudy (anonymous):

okay so what about for r(x)= SQ ROOT OVER 5-x

OpenStudy (anonymous):

\[f(x)= \sqrt{5-x}\]

OpenStudy (jwheele1):